OpenStudy (anonymous):

Hi all. I need a little help with this please. " Let one weeks supply and demand functions for gasoline be given by p=D(q)=288-5/2q and p=s(q)=3/2q, where p is the price in dollars and q is the number of 42 gallon barrels. A. Graph these equations on the same axes. B. Find the equilibrium quantity. C. Find the equilibrium price." I know no one can graph on here but if someone could show me how to obtain the values to graph it correctly, I'd appreciate it. TIA!

OpenStudy (jack1):

i would imagine you'd want to graph P (price ) vs Q ( quantity) so pick the key points of the graph: the intercept point of supply and demand, then a low point and a high point for each ie q = 1 for both and then q = 100 for both

OpenStudy (jack1):

u sure it's p = 3/2q...?

OpenStudy (jack1):

and is it: p=(288-5)/(2q) ? or \[p=288 -(\frac{ 5 }{ 2 })q ?\]

OpenStudy (anonymous):

Yes, that's it.

OpenStudy (jack1):

option 1 or option 2...?

OpenStudy (anonymous):

Actually, both. Let me double check...

OpenStudy (anonymous):

For the first one it's p=s(q)=3/2q. And that's correct for the second one.

OpenStudy (jack1):

ok, so ur 2 equations are: \[p = 288 - (\frac{ 5 }{ 2 })q \] and \[p = (\frac{ 3 }{ 2 })q\] so plug in p = 0, what are your values for q?

OpenStudy (anonymous):

All it says is q is the number of 42 gallon barrels.

OpenStudy (jack1):

no, not like that, I meant consider the equation like: if y = (3/2)x y = (1.5)x now if y = 0... what does x =...?

OpenStudy (anonymous):

Oh! Okay. I wasn't sure if I was to do the whole x=0 then y=0 thing. Okay, I got that part now.

OpenStudy (anonymous):

However, B and C are confusing me...

OpenStudy (anonymous):

Is equilibrium the same as break even?

OpenStudy (jack1):

yeah, sort of it's where the supply and demand line cross so supply price = demand price and it's only for 1 quantity value (so demand q = supply q)

OpenStudy (jack1):

 so ur supply and demand lines should look something like this when graphed

OpenStudy (jack1):

the start values you can get from subbing p=0 into the eqns and solving for q the final values u should use are if q = 0, and solve for p the other value needed is the "cross-over" value, where supply = demand quantity (q)

OpenStudy (jack1):

you can solve for this by making the equations equal eachother ie p = 3/2q = 288 - 5/2q so 3/2q = 288 - 5/2q solve for q

OpenStudy (anonymous):

At which I would add 5/2 to both sides right?

OpenStudy (anonymous):

Thank you for your help, I got it correct!

OpenStudy (jack1):

sorry i fell asleep, hey thats great, congrats